An investment offers the following yearly payments.
Year Cash Flow
1 $ 1,110
2 970
3 1,500
4 1,860
Requirement 1:

If the discount rate is 10 percent, what is the present value of these cash flows? (Enter rounded answer as directed, but do not use rounded numbers in intermediate calculations. Round your answer to 2 decimal places (e.g., 32.16).)

  Present value $   
Requirement 2:

What is the present value at 18 percent? (Enter rounded answer as directed, but do not use rounded numbers in intermediate calculations. Round your answer to 2 decimal places (e.g., 32.16).)

  Present value $   
Requirement 3:

What is the present value at 24 percent? (Enter rounded answer as directed, but do not use rounded numbers in intermediate calculations. Round your answer to 2 decimal places (e.g., 32.16).)

  Present value $   

Thank you!!!
Answer #1

Present value=Cash flows*Present value of discounting factor(rate%,time period)

a.Present value=1110/1.1+970/1.1^2+1500/1.1^3+1860/1.1^4
=$4208.12(Approx).

b.Present value=1110/1.18+970/1.18^2+1500/1.18^3+1860/1.18^4
=$3509.63(Approx).

c.Present value=1110/1.24+970/1.24^2+1500/1.24^3+1860/1.24^4
=$3099.48(Approx).